Systems Staff Meeting Notes 7/13/98

Systems Staff Meeting Notes
July 13, 1998

Present: Bill Covey, Debra Harris, Winston Harris, Sharon Henry, Suzy Shaw, Leota Whitfield

Announcements
  1. Winston & Bill came in over the weekend for the Mellon tour. Everyone was impressed with Winston's viewer, then decided they wanted to make changes. The Mellon representative was very apologetic for showing up on a Sunday.
  2. Jason Monahan will not be returning as originally determined. He has to get up to Boston earlier than anticipated for his Law School orientation.
  3. Lee sold her old Toyota Camry for a very decent price over the weekend.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Lee has received a PO that 20 Cows have arrived.
  2. A bunch of surveyed equipment was picked up today.
  3. Trudi Di Trolio's machine has 2 inches of water inside of it due to Friday's flooding. Lee will open it and see if it is wet, if so - will order new one.
  4. Room 148 - need to check on flood damage - they are still running classes in there and we don't want people getting shocked.
  5. Money has started to flow - we'll be ordering again soon - items that didn't make it in towards the end of the last fiscal year.
  6. CISCO router in MSL - Lee will get Scott Olipra to help her in retrieving it from MSL. She'll have to have the decals removed & arrange for it to be picked up.
Liaisons
  1. Need to schedule another orientation session.
  2. Need to demo Win'98 - will use the new laptop.
Networking
  1. IP management for switches in Music & Resource Services is still not working - CISCO is investigating. Suspect a network configuration problem with NERDC.
  2. Xerox public printing - every location except Education & AFA has been installed. Suspicious network slowdowns - especially in LAC - have accompanied these installations.
  3. Preservation - Should be completed this week. Workers will leave the telephone wires unterminated next to the telephone block.
  4. MSL - electrical gear is in - there was a problem in that the data conduit went to the telephone closet - it will be corrected. Furniture will probably be another six weeks.
Software Projects Status
  1. Winston met with ILL & Accounting to hear of their likes & dislikes. Basically it boils down to a "dividing of empires". He has about 4 or 5 software changes to make. Should have full migration to the SQL server by October.
  2. Mellon will be put on the back burner for now. The deadline will ultimately be at the end of the year.
  3. ATC standalone catalog needs to move up in the queue. 3270's are soon to be obsolete, need to have a backup system for a basic lookup function if the main system goes down. Winston should talk to Mark at FCLA for a copy of LUIS (our material, not all of SUS).
  4. Final Personnel Mac migration - Sharon met with Evelyn, Tom & Grace - found out that she needs to make one of their reports brand new at the start of the fiscal year and she needs to alphabetize the view to updates & deletions.
  5. Brenda Blanton (new liaison) wants to remove the Macs in Personnel - she already has an appointment with Lee to return them. They are not completely migrated yet, the permanent side for tracking has not been changed over yet.
DP Coordinator
  1. Debra will check on students returning in the Fall.
Calendar
  1. Rich will be gone to ALA/vacation June 24-July 15.
  2. Bobbie out July 13 - 17.
  3. Brian out today & July 16 & 17.
  4. Winston out July 27-31.
  5. Lee out August 17-18.
  6. Lee out Sept. 17-30


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu